The digits 1 to 7 are used to create a four digit code to disarm a bomb. How many different codes are possible if the digits may not be repeated and the code must be an odd number bigger than 5000?


1
Expert's answer
2021-01-19T04:19:33-0500

1st place: since the number should be bigger then 5000, in the 1st place we can have: 5,6 or 7 (3 possible digits)

2nd place: on the 2nd place we can put 7 – 1 = 6 digits because digits shouldn't be repeated

3rd place: on the 3rd place we can put 7 – 2 = 5 digits because digits shouldn't be repeated

4th place: odd digits we have: 1, 3, 5, 7

Therefore we have: 3 * 6 * 5 * 4 = 360 different codes are possible